can someone please help me with this question:
find the radian measure of the central angle of a circle of radius r=4inches that intercepts an arc of length s=20inches

Answer by praseena(37)

suppose 's' denotes arc length and 'r' denotes the radius of the circle,
then the ratio of any arc to the radius, s/r will be the radian measure of the central angle (theta) which that arc subtends.
So, radian measure of central angle=s/r
here s=20 inches
r=4 inches
so radian measure of central angle=20/4= 5 inches